OpenStudy (anonymous):

Solve the open sentence. –3 ≤ n + 6 ≤ 7 (Points : 1) n ≤ –9 and n ≤ 6 n ≥ –9 and n ≤ 1 n ≤ 3 and n ≤ 7 n ≥ 9 and n ≤ 7 Question 2. 2. Graph the solution of the open sentence. –5 ≤ n – 2 ≤ 2. (Points : 1) Combined Inequalities Combined Inequalities Combined Inequalities Combined Inequalities Question 3. 3. Solve the open sentence. 6 < 3b OR 3b + 12 < 9 (Points : 1) b < –1 OR b > –6 b < –1 OR b < 3 b < 2 OR b > –3 b > 2 OR b < –1

OpenStudy (anonymous):

pls answer

OpenStudy (anonymous):

okay..... add -6 to the equation -3<= n+6<=7 -3-6<=n+6-6<=7-6 -9<=n<=1

OpenStudy (anonymous):

rewrite n>=-9 , n=<1 the second one

OpenStudy (anonymous):

Q3= last one

OpenStudy (anonymous):

thank you any answers for the second question

OpenStudy (anonymous):

-3=<n=<4

OpenStudy (anonymous):